Find the slope-intercept form for the line satisfying the following conditions
x-intercept3, y-intercept 4/5

Answer by mananth(16946)

using the formula (y-y1)=m(x-x1) where (x1,y1) are the co ordinates of the point on the line
The point is (3,0)--- (x1,y1)
slope = 4/5
y-0= 4/5(x-3)
multiply by 5
5y=4(x-3)
5y=4x-12
/5
y=(4/5)x-(12/5)